A long jumper is attempting to equal the world record of 8.95 metres.

(i) Determine the angle of take-off that would enable the long jumper to attain the maximum range.

(ii) Determine the velocity of take-off that would enable the long jumper to equal the world record.

(iii) Find the maximum height attained by the long jumper while equalling the world record.

(iv)Find the time the long jumper spends in the air.


1
Expert's answer
2021-05-26T15:32:55-0400

i)


"\theta=90\degree"


ii)


"v=\sqrt{2gh}\\v=\sqrt{2(9.8)(8.95)}=13.2\frac{m}{s}"


iii)


"h=8.95\ m"

iv)


"t=\sqrt{\frac{2h}{g}}\\t=\sqrt{\frac{2(8.95)}{9.8}}=1.35\ s"